Find the integral curves of the equations
dx/(y+zx) = dy/-(x+yz) = dz/(x^2 - y^2)
1
Expert's answer
2020-06-01T17:52:39-0400

"\frac {dx}{y+zx} = \frac {dy}{-(x+yz)} = \frac {dz}{x^2 - y^2}"


"\frac {xdx+ydy}{x(y+zx)-y(x+zy)}= \frac {dz}{x^2 - y^2}"

"\frac {xdx+ydy}{z(x^2-y^2)}= \frac {dz}{x^2 - y^2}"

"xdx+ydy=zdz"

"x^2+y^2=z^2+C_1"


"\frac {xdx+ydy+zdz}{(z+1)(x^2-y^2)}= \frac {dz}{x^2 - y^2}"

"xdx+ydy+zdz=(z+1)dz"

"x^2+y^2+z^2=z^2+2z+C_2"

"x^2+y^2=2z+C_2"